These can be so confusing! Please help.
Find the product.
(8x - 1)(x^2 - 2x + 1)
(5x - 2)^2
Thank you

Answer by sdmmadam@yahoo.com(530)

1) (8x - 1)(x^2 - 2x + 1)
= 8x(x^2 - 2x + 1)-1(x^2 - 2x + 1)
(expansion by the method of first term (along with the sign) multiplied by the whole of the next bracket (addd to) second term of the first bracket(along with the sign) multiplied by the second bracket )
=8x^3-16x^2+8x-x^2+2x-1
=8x^3-16x^2-x^2+8x+2x-1
=8x^2-17x^2+10x-1 (grouping like terms)
2)(5x - 2)^2
= (5x)^2-2X(5x)X2 + 2^2
=25x^2-20x+4
[using (a-b)^2 = a^2-2ab+b^2 where a = 5x,b = 2]